### v1

# Mapping bipartite networks into multidimensional hyperbolic spaces

Abstract
Bipartite networks appear in many real-world contexts, linking entities across two distinct sets. They are often analyzed via one-mode projections, but such projections can introduce artificial correlations and inflated clustering, obscuring the true underlying structure. In this paper, we propose a geometric model for bipartite networks that leverages the high levels of bipartite four-cycles as a measure of clustering to place both node types in the same similarity space, where link probabilities decrease with distance. Additionally, we introduce B-Mercator, an algorithm that infers node positions from the bipartite structure. We evaluate its performance on diverse datasets, illustrating how the resulting embeddings improve downstream tasks such as node classification and distance-based link prediction in machine learning. These hyperbolic embeddings also enable the generation of synthetic networks with node features closely resembling real-world ones, thereby safeguarding sensitive information while allowing secure data sharing. In addition, we show how preserving bipartite structure avoids the pitfalls of projection-based techniques, yielding more accurate descriptions and better performance. Our method provides a robust framework for uncovering hidden geometry in complex bipartite systems.
I Introduction
Bipartite networks lie at the heart of countless real-world applications, linking authors to the articles they write Newman (2001a, b, c, 2004), users to the products they consume Koren et al. (2009); Konstan et al. (1997); Linden et al. (2003), people to the groups they belong to Davis et al. (2009), or countries to the languages they speak Consortium (2024). They also arise naturally in metabolic networks, where metabolites are connected to the chemical reactions or enzymes that transform them Jeong et al. (2000); Ravasz et al. (2002); Serrano et al. (2012), in plant-pollinator networks Bascompte et al. (2003); Olesen et al. (2007), and in machine learning applications, where nodes have associated features used to feed graph neural networks Aliakbarisani et al. (2023). By design, each bipartite system splits its nodes into two disjoint sets, with no edges connecting nodes within the same set. This seemingly simple rule nonetheless yields rich and complex connectivity patterns, enabling researchers to model collaboration, consumption, and association processes across diverse domains. Yet bipartite networks have historically garnered less attention than their unipartite counterparts. Given their ubiquity and straightforward interpretability, a renewed focus on bipartite structures is both timely and necessary to fully capture the multiple facets of interaction present in many real-world complex systems and data structures.
A general practice is to analyze bipartite networks by projecting them onto a single node set, creating a one-mode network Zhou et al. (2007). For instance, in an author–article bipartite network, one might create a unipartite graph of authors by connecting two authors if they have co-authored at least one article Newman (2001a, b, c, 2004). While such one-mode projections allow researchers to employ classic unipartite tools (e.g., clustering coefficients, degree distributions, community-detection algorithms), they also introduce strong correlations between edges. These correlations arise owing to sets of nodes that share common neighbors in the bipartite structure inducing cliques –or fully connected subgraphs– in the one-mode projection. Hence, the resulting unipartite networks can exhibit inflated clustering and misleading connectivity patterns that do not necessarily reflect the independent pairwise interactions of the underlying bipartite system. Besides, there is an unavoidable loss of information when bipartite networks are projected and it is even possible to obtain the same one-mode projection out of different bipartite networks Zhou et al. (2007); Neal (2014).
To overcome these limitations and more accurately capture the true structure of bipartite systems, it is critical to develop a modeling framework that treats bipartite networks directly rather than relying solely on their one-mode projections. In this paper, we propose to use network geometry to make sense of bipartite networks and to find geometric representations for this class of systems. In recent years, network geometry Boguñá et al. (2021) has been extremely successful in explaining undirected Serrano et al. (2008), directed Allard et al. (2024), weighted Allard et al. (2017), and multiplex networks Kleineberg et al. (2016, 2017) in many real systems and has also been extended to bipartite settings, proposing that nodes of both types can indeed lie in a shared latent space with connection probabilities governed by their mutual distances Serrano et al. (2012); Kitsak et al. (2017). Despite the fact that bipartite networks do not contain triangles by definition (the signature of any metric space), it is still possible to define a clustering coefficient by counting cycles of length four (i.e., squares). Empirical analyses have shown that these four-cycles can be abundant in real bipartite networks, leading to high effective clustering values Opsahl (2013); Aliakbarisani et al. (2023). This finding suggests that, indeed, bipartite networks may be embedded in a metric space where the likelihood of a connection between two nodes of different types decreases with the distance separating them.
Assuming that a bipartite network can be embedded in a similarity (or metric) space, a logical step is to seek methods for inferring the positions of its nodes within this space from real data. Previous approaches typically rely on one-mode projections of bipartite networks –using, for instance, the D-Mercator tool García-Pérez et al. (2019); Jankowski et al. (2023)– to embed each type of node. However, such projections often introduce artificial correlations and inflated clustering, reducing embedding accuracy. Here, following results in Serrano et al. (2012); Kitsak et al. (2017), we introduce the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model in which nodes of both types lie in the same $D$-dimensional similarity space. Using this model, we propose B-Mercator, an algorithm designed specifically for bipartite networks that enables the creation of multidimensional hyperbolic maps of real bipartite datasets. To illustrate its capabilities, we embeded and analyzed three datasets: Unicodelang Consortium (2024), which links countries to the languages spoken in those regions, Metabolic King et al. (2016), in which metabolites are connected through chemical reactions, and Flavor Ahn et al. (2011), which connects food ingredients to their corresponding flavor compounds. Furthermore, we show how B-Mercator can be applied to supervised machine learning tasks, including node classification and link prediction, yielding a significant performance gain with respect to state of the art methods, especially when a strong correlation exists between nodes’ labels and their feature distributions.
Figure 1: Schematic representation of the bipartite-$\mathbb{S}^{D}$ model in (a) $D=1$ and (b) $D=2$. Nodes A are shown as circles whereas nodes B are shown as squares whose sizes are proportional to the nodes’ expected degrees. The angular distances between nodes A and B are highlighted. Light grey lines represent the edges in the bipartite network generated by Eq. (1).
II Results
II.1 The bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and B-Mercator
Given a bipartite network whose links connect type-A and type-B nodes, B-Mercator finds an embedding of both node types on the surface of a $D$-sphere $\mathbb{S}^{D}$ (or, equivalently, within a $(D+1)$-sphere in the hyperbolic space $\mathbb{H}^{D+1}$), which serves as a likely realization of the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model. In the hyperbolic representation, the coordinates of nodes on the $D$-sphere represent their positions in the similarity space, whereas the radial coordinate within the hyperbolic $(D+1)$-sphere encodes their popularity, as reflected by their expected degree. Figure 1 shows two examples of bipartite networks generated by the bipartite-$\mathbb{S}^{D}$ model for $D=1$ and $D=2$. In a nutshell, pairs of type-A and -B nodes are connected with a probability that depends on their distance on the sphere, rescaled by the product of their expected degrees (represented by the size of the nodes in the figure). In the hyperbolic representation, each node is assigned a radial coordinate so that the connection probability becomes a function of the hyperbolic distance between the two nodes, thereby making the model a random geometric graph in hyperbolic geometry. In both representations, the model includes an inverse temperature $\beta_{b}$, which modulates the amount of noise in the system and thereby gauges the coupling between the network topology and its geometric characteristics. A full description of the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and the technical details of the B-Mercator embedding algorithm are provided in the Methods sections IV.1 and IV.2.
II.2 Validation
To validate our method, we generated synthetic bipartite networks from the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and measured the quality of the estimated node positions. Figure 2 shows comparisons between the true and inferred coordinates for type A and B nodes. One can observe a high Pearson correlation coefficient for $D=1$ and $D=2$, corroborating the effectiveness of our embedding technique. For more examples in $D=1$ and $D=2$ as well as $D=3$, see Supplementary Figures 2-4. B-Mercator can also infer the inverse temperature $\beta_{b}$ as shown in Supplementary Figure 5. Moreover, we tested the reproducibility of the topological properties of the original network by generating an ensemble of synthetic networks using the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model with the inferred parameters and nodes’ coordinates (see Supplementary Figures 6-10). The degree distributions and clustering spectra of type A and B nodes were very well reproduced. We also observed a good agreement between the empirical and theoretical connection probabilities. These findings confirm that B-Mercator stands out as a high-quality algorithm, accurately reconstructing the nodes’ coordinates of synthetic networks. Moreover, it reliably determines all other model parameters, such as hidden degrees and inverse temperature, regardless of the network’s dimensionality.
Figure 2: Validation of B-Mercator on synthetic bipartite networks. Relationship between the original and the inferred coordinates of the (a, b) bipartite-$\mathbb{S}^{1}$ and (c, d) bipartite-$\mathbb{S}^{2}$ models. In the top left corner of each figure, we report the value of the Pearson correlation coefficient between the inferred and original coordinates. Since the inferred coordinates for $D=2$ might be rotated, we transform them to minimize the average angular distance between the original and inferred coordinates (cf. Supplementary Section III in Jankowski et al. (2023)). Parameters for (a, b) ($D,N_{A},N_{B},\gamma_{A},\gamma_{B},\langle k_{A}\rangle,\beta_{b}$) = ($1,500,1000,2.7,2.1,10,1.5$), for (c, d) ($D,N_{A},N_{B},\gamma_{A},\gamma_{B},\langle k_{A}\rangle,\beta_{b}$) = ($2,500,1000,2.7,2.7,10,6$).
II.3 Bipartite greedy routing
In order to establish a meaningful geometric representation of a bipartite network at the global scale, we introduce bipartite greedy routing (BGR) as a practical tool to infer the network’s effective dimension Boguñá et al. (2009). The idea is to first embed the bipartite network into a latent geometric space with B-Mercator and then test how well nodes can route information by simply forwarding messages to their neighbors closest to the destination in that space. By systematically evaluating the success of these greedy routes –measured, for instance, by the probability that messages reach their targets without getting stuck– we gain insights into the dimensional structure underlying the network. A lower-dimensional latent space often requires fewer “hops” and exhibits more consistent greedy paths, thus reflecting more coherent topological structures. Conversely, if greedy routing frequently fails or requires excessive detours, it suggests a higher-dimensional or more complex geometry. In this way, the performance of greedy routing serves as an indicator of how well the bipartite graph can be embedded in a space of a given dimension, effectively allowing us to determine the dimension that best captures its structure.
We implemented a BGR protocol in which both the origin and destination can be either type A or type B nodes thus defining four variants. In Figure 3a, we depict a schematic picture of the BGR for the A–B variant, i.e., when the source node is a type A node ($S_{A}$), and the target node is a type B node ($T_{B}$). The message is forwarded from $S_{A}$ to the type B node that is hyperbolically closest to the target. Since the node $B_{1}$ is not the destination, the message is forwarded again to a type A node. The process is repeated until the destination is reached or the message becomes stuck. Then, the BGR protocol is executed for a large number of randomly chosen node pairs to assess the global network’s geometric properties.
We tested the BGR protocol in synthetic networks generated from the bipartite-$\mathbb{S}^{D}$ model. First, we generated networks with specific dimensionality and topological properties, and we obtained their hyperbolic maps by embedding them using B-Mercator with different embedding dimensions. The performance of BGR was assessed based on two key measures: the proportion of messages that successfully reach their destination $p_{s}$, i.e., the success rate, and the mean stretch, where stretch is defined for each path connecting a source and target node as the ratio of the hop count of a successful greedy path to the shortest path. In Figure 3b, we show the success rate as a function of the embedding dimension for networks generated using the bipartite-$\mathbb{S}^{D}$ model with dimensions ranging from $D=1$ to $D=4$. Interestingly, the performance of BGR is optimal (in terms of $p_{s}$ and mean stretch, see Supplementary Figure 12) when B-Mercator is used with the same dimension that was used to generate the network, thus justifying BGR as an alternative method to infer the effective dimension of real networks, different from the topological-based method introduced in Almagro et al. (2022). These results are consistent across all variants of BGR (see Supplementary Figure 11) and corroborate the results obtained for unipartite networks Jankowski et al. (2023).
Figure 3: Bipartite greedy routing in synthetic networks. (a) Schematic view of the greedy routing protocol. We select a type A node as an origin ($\mathbf{S_{A}}$) and a type B ($\mathbf{T_{B}}$) as a destination. The red arrows show how the package/message is forwarded towards the destination. In the second example, we select two type B nodes as source ($\mathbf{S_{B^{\prime}}}$) and destination ($\mathbf{T_{B^{\prime\prime}}}$) and outline the greedy path with purple color. The line width is proportional to the connection probability (Eq. 1). (b) Success rate as a function of embedded dimension for networks generated in $D=\{1,2,3,4\}$. We consider here a navigation protocol where a source is a type A node and a destination a type B one. Results are obtained by averaging over 10 realizations with $(N_{A},N_{B},\gamma_{A},\gamma_{B},\left<k\right>,\beta_{b})$ = $(500,500,2.5,2.5,10,1.5)$.
II.4 Embedding of real bipartite networks
The significance of B-Mercator lies not only in its ability to embed synthetic networks generated by the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model, but rather in its capacity to uncover geometric insights from real bipartite networks. Moreover, embeddings produced by B-Mercator can be applied to tasks such as node classification and link prediction on graph-structured data. As case studies, we analyze the Unicodelang dataset, which captures relationships between countries and the languages spoken within them, the human metabolic network connecting metabolites with the reactions in which they participate, and the flavor network linking ingredients to the chemical compounds they contain. These examples demonstrate the practical applicability of B-Mercator in extracting meaningful structural patterns from real-world bipartite networks.
Using B-Mercator, we embedded the Unicodelang dataset in various dimensions. The inferred embeddings are able to reproduce the topological properties of the network (see Supplementary Figure 15). In Figure 4, we show a dual embedding representation of this dataset in dimension $D=1$. First, we focus on the countries in which a given language is used. For instance, in Figure 4a, we plot all countries in which English is spoken, i.e., the neighbors of the English language in the bipartite network. One can notice that English is located closer to India or the Philippines than to the United States of America (USA). This can be explained by the fact that many different languages are spoken in the USA, which influence its position in the bipartite map. Indeed, in Figure 4b, where we plot all neighbors of the Spanish language, the USA is located close to Spanish language. Lastly, in Figure 4c one can observe that countries from the French colonialism are concentrated in the similarity space and lie close to the French language, in contrast to France and other European countries. We can also shift the perspective: instead of examining the neighbors of each language, we can analyze the neighbors of a given country. In the bottom row of Figure 4, we plot the language neighbors of India, China, and Brazil. Countries are often located in the embedding space close to the most widely spoken language. Additional examples are provided in Supplementary Figure 13, where we explore the neighbors of languages such as Hindi, Swahili, Catalan, Persian, Korean, Dutch, Russian, and Arabic. Similarly, in Supplementary Figure 14, we depict the neighbors of countries including Canada, Turkey, Indonesia, Cameroon, Tanzania, Greece, the Philippines, and Bolivia.
Figure 4: Visualization of the bipartite-$\mathbb{S}^{1}$ embedding of the Unicodelang dataset per country or language. Panels (a, b, c) show countries where a given language is spoken, i.e., the neighbours of the language node. The size of the nodes is proportional to the number of language speakers in that country. The color corresponds to the geographical region in which the country is located. A star marker indicates the position of a given language. In panels (d, e, f), we depict all languages spoken in a given country, i.e., the neighbours of the country node. The size of the nodes is proportional to the fraction of speakers of a given language. The color represents that language’s script. A cross marker indicates the position of a given country.
The hyperbolic bipartite embedding enables us to examine the concentration of countries sharing the same language. We select the top 15 languages with the highest degree and compute the angular distance from each language to its neighboring countries. A small average angular distance may indicate that these countries are highly similar, whereas a broad distribution suggests a more international language. Figure 5a displays the diversity for these high-degree languages. As expected, the angular distance distribution for English is broad, indicating connections with countries distributed throughout the similarity space. Interestingly, the angular distances for Fula are relatively small. Fula, a Senegambian language spoken primarily in West and Central Africa, is concentrated in a specific region of our embedding space. Similarly, we can examine the linguistic diversity within each country. Following the previous approach, we selected the top 15 countries with the highest degree and computed the angular distances to each neighboring language. In a given country, if the languages are more similar, their angular positions in the similarity space should be more concentrated. Figure 5b shows that India and Canada exhibit a broader distribution of angular distances, reflecting the presence of a diverse array of languages. In contrast, Russia and Cameroon display a narrower distribution, indicating a more homogeneous set of languages. Notably, all these countries are linguistic hubs, with approximately 30 or more languages spoken in each. These results suggest that our embedding can serve as an indicator of a language’s international reach and linguistic diversity, which is not solely reflected by its degree.
Figure 5: Language or country diversity for the top 15 highest-degree nodes. Panel (a) shows violin plots of angular distances between each language and its neighboring countries, with colors indicating script type. Panel (b) presents analogous plots for countries, with colors representing geographic region. In both panels, nodes are ordered in descending order of degree, with each node’s degree (in brackets) indicated next to its label. Only the central 95% of the data is plotted—that is, data between the 2.5th and 97.5th quantiles are shown. A black line highlights the median value in each plot.
In addition, we investigated the human metabolic network, defined as metabolites connected to the reactions they participate in King et al. (2016) and the network of food ingredients based on the flavor compounds they share Ahn et al. (2011) (see Section IV.3). In both cases, B-Mercator is able to reproduce topological properties of this biparite network such as the degree distributions and clustering spectra. See Supplementary Section 6 for more details.
Finally, we applied the bipartite greedy routing protocol to the embeddings derived from these real-world networks. For the Unicodelang network, the highest success rate—based on the $A-A$ BGR variant—is observed at an embedding dimension of $D=4$. However, variations in the success probability ($p_{s}$) across different embedding dimensions are minimal, likely due to the relatively low value of $\beta_{b}$. In contrast, the highest $p_{s}$ for both the Metabolic and Flavor networks is achieved at $D=1$. A summary of these findings is provided in Supplementary Table 2.
II.5 Case study on the graph machine learning tasks
Graph Machine Learning (Graph ML) focuses on extracting patterns, making predictions, and uncovering insights from graph-structured data Hu et al. (2020); Xia et al. (2021). This data is typically defined as a set of entities (nodes) with complex relationships (links), defining a unipartite graph $\mathcal{G}_{n}$. Nodes are enriched with a set of features, defining a bipartite network of nodes and features $\mathcal{G}_{n,f}$. Within the Graph ML community, two key tasks are commonly used to evaluate and rank network embedding models: node classification (NC) and link prediction (LP). In turn, network embeddings can be broadly categorized into supervised and unsupervised approaches. Supervised embeddings, such as those learned by graph neural networks (GNNs), use node labels in the training set to inform the learning process for classification tasks. In contrast, unsupervised methods leverage only the network structure and, optionally, node features to generate low-dimensional representations of the data. These maps can then be used for multiple downstream tasks by integrating additional classification models.
Our embedding method B-Mercator belongs to the class of unsupervised graph embeddings that leverage only the nodes’ feature matrices. This is possible thanks to the findings in Aliakbarisani et al. (2023), which show that nodes and their associated features define a bipartite network, $\mathcal{G}_{n,f}$, with strong geometric properties. We used B-Mercator to find an embedding of the nodes-features bipartite network in the common similarity space, which we subsequently used to perform NC and LP tasks in a supervised manner. To highlight the importance of such embeddings, we compared B-Mercator with D-Mercator Jankowski et al. (2023), which produces multidimensional hyperbolic maps of unipartite networks $\mathcal{G}_{n}$ without using information from the nodes’ features. We selected embedding dimensions $D=1$ and $D=2$ to map the node features into the bipartite-$\mathbb{S}^{1}$ and bipartite-$\mathbb{S}^{2}$ models using B-Mercator, and to map the unipartite network into the $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$ models using D-Mercator.
We also compared our model-driven approach with existing state-of-the-art graph embedding methods in both node classification and link prediction tasks. For further details on the methods used, see Supplementary Section 7. We selected seven graph datasets commonly used in machine learning research, each with varying levels of correlations between the graph, nodes’ features, and nodes’ labels, (see Supplementary Section 8 for details). It has been shown recently that the performance metrics of Graph ML tasks can vary significantly depending on these correlations Jankowski et al. (2024). For instance, in the node classification tasks, adding features can be detrimental when the correlation between node features and network structure is very low. In contrast, adding features significantly enhances the results when the correlation is high. Among the analyzed networks, Cora and Citeseer exhibit a strong correlation between network structure and node features. In contrast, IMDB, Wisconsin, Texas, and Cornell show relatively low correlation, while Film demonstrates almost no correlation (see Supplementary Table S3 for more details).
Figure 6: Case study on machine learning tasks. (a) Accuracy of the node classification task and (b) balanced precision of the distance-based link prediction task for the Wisconsin dataset. The train/test split is 20/80 for nodes in NC and 90/10 for links in LP, where the test set is balanced by randomly adding an equal number of negative links selected from non-existing links. The results are averaged over 10 different splits. Our methods are highlighted with diagonal hatches. The abbreviations of the algorithms are as follows: DW – DeepWalk, R2V – Role2Vec, LE – Laplacian Eigenmaps, FN – FeatherNode. The numeric value in brackets indicates the embedding dimension. All other parameters are set to their default values. The methods are sorted by the median accuracy in NC and balanced precision in LP tasks. In addition, the methods are grouped and colored by input data type, i.e., Network-based methods use only network topology, Feature-based methods use only nodes’ feature matrix, and Feature$+$Network-based methods merge two things to construct the network embedding. (c) Visualization of the bipartite embedding with B-Mercator in $D=1$ for the Wisconsin dataset. We plot the positions of the nodes and color them based on the metadata. (d) Average rank in link prediction and node classification tasks across seven datasets. We plot the medium rank value with the error bars depicting the interquartile range (IQR). The size of the markers is inversely proportional to the embedding dimension and their shapes correspond to the input data type. Our methods are outlined with a thicker marker border.
To compute the accuracy of the node classification task, we applied a KNeighborsClassifier from the scikit-learn library Pedregosa et al. (2011) to each network embedding with $K=10$ as the number of nearest neighbors. We split the data into training and testing subsets with a 20/80% ratio. In the case of B-Mercator and D-Mercator, we computed distances among pairs of nodes using their angular separation on the $D$-sphere. For the rest of the methods, we computed the Euclidean distance between the nodes’ positions from the corresponding embedding. In Fig. 6a, we report the performance of the NC task on the Wisconsin dataset. This is a network of web pages from the Computer Science department of the University of Wisconsin. Each web page is enriched with 1613 features and manually classified into one of five categories: student, project, course, staff, and faculty. For this dataset, feature-based methods demonstrate superior performance, with our methods (B-Mercator in $D=1$ and $D=2$) achieving a significant margin of improvement over competitors. This is due to the strong correlation between the nodes’ features and labels, as shown in Jankowski et al. (2024). In contrast, network-based methods tend to perform less effectively, as they rely solely on the graph topology, which is weakly correlated with the nodes’ labels Jankowski et al. (2024).
For the distance-based link prediction task, we randomly selected a small fraction $q=0.1$ of the existing links from the network structure as positive links. Similarly, an equal number of non-existing edges was selected as negative links to form a balanced test set. The remaining positive links were considered as the training network, which was then embedded using various network-based methods. In the case of B-Mercator, the bipartite network between nodes and features was embedded, and it was not affected by splitting the links in the network structure into training and test sets. For B-Mercator and D-Mercator, we used the inverse of the hyperbolic distance between node pairs as the similarity measure, while for the other methods, we used the inverse of the Euclidean distance between node positions. As a result, links in the test set with the smallest distance were ranked highest when sorted in ascending order of their similarity scores.
In Fig. 6b, we show the precision of the different methods. In this case, methods using only the network topology, or a combination of network structure and node features, deliver the highest precisions. For this task, our embeddings with D-Mercator ($D=1$ and $D=2$), using only the network topology, are competitive and achieve a performance similar to MUSAE in dimension $D=8$, demonstrating the versatility of low-dimensional embeddings derived from our approach. Moreover, among feature-based methods, B-Mercator in $D=1$ and $D=2$ outperforms all other approaches. This highlights the adaptability of our methods to different tasks depending on the underlying data representation and task requirements. Detailed results for additional datasets can be found in Supplementary Figures S23–S46.
Figure 6d provides a summary of the results across multiple datasets. For each dataset, the rank of each method is calculated for both NC and LP tasks, and the median rank is plotted along with the interquartile range (IQR). B-Mercator in low dimensions ($D=1$ and $D=2$) consistently ranks among the top approaches for the NC task and outperforms all feature-based methods for the LP task, for which D-Mercator is the best method in dimension $D=1$. These results demonstrate the reliability and adaptability of our methods across datasets. The marker size in the plot, which inversely represents the embedding dimension, further underscores the efficiency of our low-dimensional embeddings compared to high-dimensional alternatives. In Supplementary Figures S30 and S38, we provide a more detailed view of the rankings for the NC and LP tasks.
III Discussion
The ability to embed real-world systems into a geometric space is a pivotal step toward understanding their intrinsic structure, function, and underlying organization. While numerous network embedding techniques have been proposed, most are not derived from a model-based perspective, limiting their interpretability and capacity to reconstruct empirical data. Model-based approaches offer a principled way of capturing the generative mechanisms that shape network topologies, thereby enabling researchers to interpret embeddings in a manner closely aligned with the data’s underlying structure.
Several model-based embedding methods have already proved effective for unipartite networks, yet comparable solutions for bipartite networks have remained underexplored. In this work, we addressed this gap by introducing B-Mercator, a novel geometric model-based embedding algorithm specifically designed for bipartite networks. By mapping bipartite structures into hyperbolic space, B-Mercator offers a powerful and interpretable way to capture community structure, hierarchical organization, and topological relationships.
To demonstrate the versatility of B-Mercator, we applied it to embed real-world bipartite systems. The analysis of the language network (countries–spoken languages), the metabolic network (metabolites–reactions), and the Flavor network (ingredients–chemical compounds) show that the embeddings not only correlate well with metadata but also retain the essential characteristics of each dataset. Furthermore, we evaluated B-Mercator on several node classification and link prediction tasks. It consistently outperformed all unsupervised methods for node classification and emerged as the best performer among feature-based embeddings for link prediction.
We stress an important advantage of our model-based embeddings. The hyperbolic maps can be used to generate synthetic networks with node features that closely resemble their real-world counterparts (see Supplementary Figures 46-48). By doing so, we safeguard any sensitive information derived from real complex networks, such as personal connections, transactional data, or proprietary interactions. Thus, we enable the secure sharing of structural data without compromising the integrity of the original network or revealing sensitive information.
These findings underscore the value of geometric model-based embeddings for both theoretical analyses and practical applications, ranging from community detection and studying network hierarchies to advanced machine learning tasks. B-Mercator’s robust performance highlights its capacity to reveal meaningful insights into bipartite systems –a domain often overlooked in the current embedding literature– while providing substantially more accurate analyses than those based on corresponding one-mode projections. Overall, B-Mercator represents a significant advancement in bipartite network analysis, paving the way for more accurate, interpretable, and generative representations of complex real-world systems.
IV Methods
IV.1 Bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model
In the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model – an extension of the bipartite-$\mathbb{S}^{1}/\mathbb{H}^{2}$ Aliakbarisani et al. (2023) – we assign to each node (of type A or B) a hidden degree ($\kappa_{A}$ or $\kappa_{B}$) and the position in the $D$-dimensional similarity space chosen uniformly at random, and represented as a point on a $D$-dimensional sphere. Each node of type A and B is assigned a vector $\mathbf{x}_{i}\in\mathbb{R}^{D+1}$ with $||\mathbf{x}_{i}||=R$. For instance, when $D=1$ the similarity space is represented as a circle, whereas for $D=2$ it is a sphere (see Figure 1).
The connection probability between node $u$ of type A and node $v$ of type B takes the form of gravity law:
${p}_{uv}=\frac{1}{1+{\chi}_{uv}^{\beta_{b}}},\,\,\,\mathrm{with}\,\,\,{\chi}_{uv}=\frac{R{{\Delta}}{\theta}_{uv}}{{\left(\mu_{b}{\kappa}_{u}{\kappa}_{v}\right)}^{1/D}}.$ (1)
The number of nodes of type A (type B) is $N_{A}$($N_{B}$), for convenience and without loss of generality, we set the density of nodes of type A in the $D$-sphere to one so that
$R={\left[\frac{N_{A}}{2{\pi}^{\frac{D+1}{2}}}{{\Gamma}}\left(\frac{D+1}{2}\right)\right]}^{\frac{1}{D}}.$ (2)
The separation $\Delta\theta_{uv}=\arccos\left(\frac{\mathbf{x}_{u}.\mathbf{x}_{v}}{R^{2}}\right)$ represents the angular distance between nodes $u$ and $v$ in the $D$-dimensional similarity space. The parameter $\beta_{b}$ (with $\beta_{b}>D$) controls the coupling between the resulting topology and the underlying metric space. Lastly, the parameter $\mu_{b}$ controls the average degree of nodes of type A and is defined as
$\mu_{b}=\frac{\beta_{b}{{\Gamma}}\left(\frac{D}{2}\right)\sin\frac{D\pi}{\beta_{b}}}{2{\pi}^{1+\frac{D}{2}}\langle k_{A}\rangle},$ (3)
whereas the average degree of type B nodes is set by $\langle k_{B}\rangle=\frac{N_{A}}{N_{B}}\langle k_{A}\rangle$. By choosing distributions for the hidden degrees $\kappa_{A}$ and $\kappa_{B}$ and inverse temperature $\beta_{b}$, we can generate bipartite networks with any desired degree distributions and varying levels of geometric properties.
Interestingly, the bipartite-$\mathbb{S}^{D}$ model can be represented in purely geometric terms as the bipartite-$\mathbb{H}^{D+1}$ model. This is achieved by mapping the hidden degrees of each type A and B nodes to radial coordinates while preserving their positions on the sphere $\mathbb{S}^{D}$. Specifically, the transformation for type A nodes has the form (similarly for type B nodes)
${r}_{u}=\hat{R}-\frac{2}{D}\ln\frac{{\kappa}_{u}}{{\kappa}_{u,0}},\,\,\,{{\rm{with}}}\,\,\,\hat{R}=2\ln\left(\frac{2R}{{(\mu_{b}{\kappa}_{u,0}{\kappa}_{v,0})}^{1/D}}\right).$ (4)
where ${\kappa}_{u,0}$(${\kappa}_{v,0}$) is the smallest hidden degree for type A (type B) nodes. We can rewrite Eq. (1) as
${p}_{uv}=\frac{1}{1+{e}^{\frac{\beta_{b}}{2}({x}_{uv}-\hat{R})}},\,\,\,{{{{{{{\rm{with}}}}}}}}\,\,\,{x}_{uv}={r}_{u}+{r}_{v}+2\ln\frac{{{\Delta}}{\theta}_{uv}}{2}.$ (5)
With this transformation, the space represented by the radial position of each node, along with its angular position on the sphere, becomes the hyperboloid model of the hyperbolic space of dimension $D+1$. Consequently, the connection probability between nodes $u$ and $v$ becomes a function of $x_{uv}$, which is a good approximation of the hyperbolic distance between them.
IV.2 B-Mercator in details
We adopt the code of $D$-Mercator Jankowski et al. (2023) for the embedding bipartite networks. Here, we provide the overview of the differences between the embeddings for the unipartite and bipartite networks.
Inferring the hidden degrees and parameter $\beta_{b}$.
The inference of hidden degrees for type A and B nodes, and the inverse temperature $\beta_{b}$ is implemented as an iterative process. We begin with the initial guess for the parameter $\beta_{b}\in(D,2D)$, where $D$ is an embedding dimension, and initialize the hidden degrees as the observed degrees in the original network for nodes A ($\{k_{A,i},i,\dots,N_{A}\}$) and nodes B ($\{k_{B,i},i,\dots,N_{B}\}$). The aim of the estimation is to modify the hidden degrees in order to ensure that the expected degree of each node within the model aligns with the degree observed in the original network. After the hidden degrees for both nodes A and B are computed, the synthetic graph from bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ is constructed and the bipartite clustering coefficient is calculated. If the computed bipartite clustering coefficient deviates from that of the original network, $\bar{c}_{b}$, the value of $\beta_{b}$ is adjusted. Then, the process is repeated using the current estimation of hidden degrees until a predetermined precision is reached.
bipartite-$\mathbb{S}^{D}$ model corrected Laplacian Eigenmaps.
Since the biadjacency matrix of a bipartite graph $\mathbf{A}$ is not symmetric, to apply the Laplacian Eigenmaps, we transform it to the adjacency matrix as
$\mathbf{B}=\begin{bmatrix}0&\mathbf{A}\\ \mathbf{A}^{T}&0\end{bmatrix}$ (6)
Similarly to Jankowski et al. (2023), the expected angular distance between nodes $u$ (of type A) and $v$ (of type B) in the bipartite-$\mathbb{S}^{D}$ model, conditioned to the fact that they are connnected, can be computed as
[…]
Additionally, for $D=1$, we keep the ordering inferred by LE and distribute nodes of type A and B evenly on the circle.
Likelihood maximization
The nodes’ coordinates in the similarity space inferred using LE are adjusted by Maximum Likelihood Estimation (MLE) to optimize the probability that the bipartite-$\mathbb{S}^{D}$ model generates the observed network. We define an order of nodes sorted by their degree for A and B type nodes separately. Fixing the positions of a subset of B type nodes, we find new optimal coordinates for a subset of type A nodes. First, we compute the mean coordinates of type A node $u$’s neighbors.
$\mathbf{x}_{u}=\sum\limits_{v}\frac{1}{\kappa_{v}^{2}}\mathbf{x}_{v}$ (8)
where the sum goes of all neighbors of node $u$, i.e., nodes of type B. Later, the new positions around $\mathbf{x}_{u}$ are proposed using a multivariate normal distribution. Finally, we select the most likely candidate position based on the local log-likelihood
$\ln\mathcal{L}_{u}=\sum\limits_{v=0}^{N_{B}}a_{uv}\ln p_{uv}+(1-a_{uv})\ln(1-p_{uv})$ (9)
After iterating over a subset of nodes of type A, we apply a similar approach to nodes of type B, and repeat the process until all node positions are adjusted.
Final adjustment of hidden degrees
Lastly, we adjust the hidden degree to compensate deviations from $\bar{k}_{A}(\kappa_{u})=\kappa_{u}$ and $\bar{k}_{B}(\kappa_{v})=\kappa_{v}$, which might have been introduced during the estimation of the nodes’ coordinates in the similarity space.
IV.3 Real bipartite datasets
The Unicodelang dataset was downloaded from the Unicode CLDR Project GitHub repository Consortium (2024). The dataset [^1] https://github.com/unicode-org/cldr/blob/main/common/supplemental/supplementalData.xml contains information about the languages spoken in a given territory. We preprocessed this dataset and matched the country codes to their geographical regions [^2] https://github.com/lukes/ISO-3166-Countries-with-Regional-Codes. In total, the bipartite graph contains 246 countries, 717 languages, and 1487 edges.
The metabolic network was extracted from the BiGG webpage King et al. (2016). We focus on the RECON1 model, which corresponds to Homo sapiens. The data was preprocessed and cleaned, resulting in 1497 metabolites and 2212 chemical reactions.
The flavor network is a network of food ingredients based on the flavor compounds they share Ahn et al. (2011). After preprocessing, the total number of ingredients is 602, and the number of compounds is 1138. The bipartite graph contains 15,382 edges. Additional network properties and inferred values of $\beta_{b}$ for different embedding dimensions are shown in Supplementary Table 1.
Acknowledgments
We acknowledge support from: Grant TED2021-129791B-I00 funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR”; Grant PID2019-106290GB-C22 funded by MCIN/AEI/10.13039/501100011033; Generalitat de Catalunya grant number 2021SGR00856. R. J. acknowledges support from the fellowship FI-SDUR funded by Generalitat de Catalunya. M. B. acknowledges the ICREA Academia award, funded by the Generalitat de Catalunya.
V Data availability
The network datasets used in this study are available from the sources referenced in the manuscript and the Supplementary Information.
VI Code availability
The open-source code for B-Mercator, along with the code to reproduce the figures, will available on GitHub at https://github.com/networkgeometry/b-mercator upon publication.